SITE 1038 (PROPOSED SITE ET1-4)
The ship was moved in DP mode 3.1 nmi in 4.5 hr at 0.7 kt to Site 1038 at dGPS coordinates 41°00.012'N, 127°29.633'W. Datasonics beacon S/N 1247 was dropped at 1556 hr on 2 October. Operations were shut down for 1 hr to reinstall and test the DCS sensor electronic equipment. The VIT-TV was run, and a TV survey was conducted over Site 1038 for 0.75 hr. Marker 6X (deployed by submersible) was located, and Hole 1038A (41°00.0163'N, 127°29.6851'W) was spudded at 0130 hr on 3 October in sulfides. The seafloor was tagged at 3348.0 mbrf. The TV was pulled, and RCB Cores 169-1038A-1R to 12R were taken from 0 to 114.5 mbsf (3348.0 mbrf) with 2.3% recovery. Core 169-1038A-1R encountered massive sulfides, but recovery was very poor in the remaining cores. The TV was run to check the hole and confirmed that there was no flow from Hole 1038A. The hole at the seafloor was ruggose and about 18 in. wide. The bit cleared the seafloor at 2030 hr on 3 October.
The ship was moved in DP mode 300 m west of Hole 1038A, and the bottom was checked with the TV. Hole 1038B (41°00.0247'N, 127°29.8707'W) was spudded at 2240 hr on 3 October. The seafloor was tagged at 3266.0 mbrf. RCB Cores 169-1038B-1R to 13R were taken from 0 to 120.5 mbsf (3386.5 mbrf) with 11.5% recovery. The bit cleared the seafloor at 1700 hr on 4 October. The TV was not run until after pulling out of the hole because there were three crossing swells from remote storms, which caused rolls to 4 and heave to 2.75 m, and the ship could not be rotated because the current was shifting from 0 to 4 kt and abruptly changing direction by more than 90.
The ship was moved in DP mode 76 m east and 259 m north of Hole 1038A. A site was selected on top of the mound on the north edge with the TV survey. Hole 1038C (41°00.1816'N, 127°29.6037'W) was spudded at 2200 hr on 4 October. The seafloor was tagged at 3239.0 mbrf. RCB Cores 169-1038C-1R to 4R were taken from 0 to 41.8 mbsf (3280.8 mbrf) with 7.6% recovery. The bit cleared the seafloor at 0515 hr on 5 October.
The ship was moved in DP mode 12 m west and 64 m south of Hole 1038A. A site was selected with the TV beside a pile of what appeared to be sulfide chimney fragments and boulders. The seafloor was tagged at 3237.0 mbrf, and Hole 1038D (41°00.1441'N, 127°29.6146'W) was spudded at 0745 hr on 5 October. RCB Cores 169-1038D-1R to 4R were taken from 0 to 43.8 mbsf (3280.8 mbrf) with 7.4% recovery. The bit cleared the seafloor at 1540 hr on 5 October.
The ship was moved 300 m south to Hole 1035A, which still had no flow coming from it. The ship was moved about 35 m east and 31 m north of Hole 1038A. The seafloor was tagged at 3221.0 mbrf, and Hole 1038E (41° 00.0012'N, 127°29.6463'W) was spudded at 1845 hr on 5 October. RCB Cores 169-1038E-1R to 4R were taken from 0 to 40.6 mbsf (3261.6 mbrf) with 5.9% recovery. The bit cleared the seafloor at 0015 hr on 6 October.
The ship was moved 435 m south and 525 m east of Hole 1035A to the edge of a sulfide mound. The seafloor was tagged at 3252.0 mbrf, and Hole 1038F (40°59.7304'N, 127°29.2877'W) was spudded at 0615 hr on 6 October. RCB Cores 169-1038F-1R to 4R were taken from 0 to 38.4 mbsf (3290.4 mbrf) with 17.2% recovery. The RCB BHA was pulled, and the bit cleared the rotary at 1745 hr on 6 October.
We made a pipe trip and the standard APC/XCB BHA was run to obtain sediment cores near the sulfide mound outcrops on the east side of Bent Hill. The plan was to move the ship 5 m west of Hole 1038F to a 3- to 4-m-deep and 20-m-wide depression in the sulfide deposit. Clams in the depression indicated diffuse flow and the seafloor was tagged at 3251.0 mbrf, and Hole 1038G (40°59.7255'N, 127°29.2865'W) was spudded at 0245 hr on 6 October. XCB Cores 169-1038G-1X to 2X were taken from 0 to 22.0 mbsf with no recovery to penetrate through potentially hard sulfides just below the seafloor. The bit penetrated through a firm upper surface on Core 1X into soft sediments; therefore, APC coring was initiated. APC Cores 169-1038G-3H to 5H were taken from 22.0 to 50.5 mbsf. Cores 169-1038G-3H and 4H were noticeably warmer and very gassy, and the core liners were blackened by heat. On Core 169-1038G-5H the liner was softened and pushed back up into the liner by the core, which jammed the liner and doubled the thickness of the liner tube (like a tubing upset). XCB Cores 169-1038G-6X to 16X were taken from 50.5 to 147.0 mbsf (3398.0 mbrf). Basalt was encountered on Cores 169-1038G-15X and 16X. Coring was terminated when the XCB shoe was destroyed on Core 169-1038G-16X.
The ship was moved in DP mode back to Hole 1038A at 41°00.016'N, 127° 29.685'W. A TV survey was conducted. The seafloor was tagged at 3234.0 mbrf, and Hole 1038H was spudded at 1030 hr on 8 October. XCB Cores 169-1038H-1X to 20X were taken from 0 to 192.8 mbsf with 16% recovery. The bit cleared the sea floor at 1400 hr on 9 October.
The ship was moved 250 m south and 130 m east from Hole 1038A onto Central Hill. An XCB core barrel was run. The seafloor was tagged with the bit at 3227.0 mbrf. Hole 1038I was spudded at 1600 hr on 9 October. XCB Cores 169-1038I-1X to 2X were taken from 0 to 17.3 mbsf with 55.7% recovery. APC Cores 169-1038I-3H to 7H were taken from 17.3 to 64.8 mbsf with 102.0% recovery. An Adara temperature measurement was taken on Core 169-1038I-3H and the Davis-Villinger temperature probe (DVTP) was run after Core 169-1038I-6H at 55.3 mbsf. The last three APC cores were partial strokes with 50-60K lb overpull so APC coring was terminated. XCB Cores 169-1038I-8X to 43X were taken from 64.8 to 404.0 mbsf (3631.0 mbrf) with 40.0% recovery. Coring was terminated when time for operations on Leg 169 expired.
Coring was finished at Hole 1038I, and about 123 stands (1200 m) of 5" drill pipe had been pulled to prepare for the transit to Los Angeles. The seafloor had been cleared, and the ship had been moved slowly in DP mode about 0.4 nmi southeast of Central Hill (in preparation for changing out the coring wire line) to 40°59.60'N, 127°28.98'W. While pulling pipe at about 1700 hr, the roughnecks attempted to put both bails under the pickup elevator ears, but only one bail engaged. When the elevators were picked up, the engaged bail tilted the elevator, slipped off the ear, and broke off the end of the elevator-ear latch. The elevators flipped up on one side, and the 5" drill pipe bent and broke 0.7 m below the top of the tool joint box. No one was injured. The incident can be attributed to combined operator and crew error, with end of leg fatigue as an extenuating factor. Lost on the seafloor were: 81 stands (2298 m) of 5" drill pipe, and 127 m of BHA (crossover, 2 stands of 5-1/2" drill pipe, tapered drill collar, 5 X 8-1/4" drill collars, HS, TS, LS, SBDC, bit sub, 11-7/16" Sec S86F bit).
To 169 Operations Summary
To 169 Table of Contents